# Classification of foliations via disintegration maps

### v1

@@ 3. Preliminaries on disintegration maps
In this section, we discuss key properties of disintegration maps, as introduced in Definition 2.1 and developed in [PR25]. To start our discussion, suppose $X=Y=\mathbb{R}$, and let $\gamma$ be a probability measure supported on the graph of a measurable function $\pi:X\to Y$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ its a probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the pushforward of $\mu$ under the map $x\mapsto(x,\pi(x))$. Then, the disintegration of $\gamma$ with respect to $\mu$ is given by the family of Dirac measures:
$\gamma_{x}=\delta_{\pi(x)}.$
In fact, for any measurable $A\subseteq X\times Y$, we have
@@
By definition of the disintegration
$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x).$
where $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
@@
In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration — specifically, through the properties of disintegration maps. A key starting point for this analysis is understanding the continuity of these maps. When referring to continuity of the disintegration map, we mean continuity with respect to weak convergence on $\mathscr{P}_{p}(X)$. In [PR25] some results are presented regarding the continuity of disintegration maps, namely:
@@
    [PR25, Propostion 5.2] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If $\pi:X\to Y$ is a Borel map such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$. A map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if, for each $\varepsilon>0$ there exist $\mathcal{K}\subset Y$ closed with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
@@
    [PR25, Propostion 5.4] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If the Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
@@
    [PR25, Propostion 5.8] Let $X$ be locally compact complete separable metric space. Considering a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$, if there exists a metric measure foliation of $X$, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
The continuity of disintegration maps enables us to relate paths in the chosen indexing metric space to corresponding paths of measures in Wasserstein space. For instance, in the case of the disintegration map $x\mapsto\delta_{x}$, there is a direct correspondence between a path in $X$ and the induced path of Dirac measures. In the context of a metric measure foliation, this relationship becomes more nuanced. Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path — a consequence of the parallelism of the leaves.
In the next section we present an approach that allows us to study the relation between the relative position of the elements of the family of supports and the relative position of the conditional measures in the Wasserstein space, via a disintegration map.
@@ 4. On foliations and disintegration maps
Motivated by the concept of the metric derivative and the interplay between the relative positions of supports and conditional measures, we introduce a notion of derivative for disintegration maps to quantify how these measures evolve along paths.
@@
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, $\pi:X\to Y$ a Borel map, $\nu=\pi_{*}\mu$, and $f$ the disintegration map of $\mu$ with respect to $\nu$. We define a notion of derivative of $f$ by
$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}\sup_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))}$
for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures and $B_{\varepsilon}(y)$ is a ball centered in $y\in Y$ with radius $\varepsilon$.
Let $d_{Y}(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$ denote the distance between the fibers over $y^{\prime}$ and $y^{\prime\prime}$. In the case of a metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d_{Y}(y^{\prime},y^{\prime\prime})$, which implies $|\nabla f(y)|_{p}=1$. This derivative concept thus serves as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a metric measure foliation. To further analyze the connection between the distances of conditional measure supports and their Wasserstein distance, paralleling techniques in geometric analysis, we introduce a dense energy functional designed to quantify this relationship.
Definition 4.2.
Let $X,Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$, a Borel map $\pi:X\to Y$, and the disintegration map $f$ of $\mu$ with respect to $\nu=\pi_{*}\mu$. The $p$-energy of $f$ is given by
Note that $|\nabla f|_{p}\geq 1$ everywhere. In fact, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
@@
Moreover, note that $\min_{f}\mathcal{E}_{p}(f)=1$ if, and only if, $\|\nabla f\|_{\infty,p}=1$, which in turn hold true if, and only if $|\nabla f|_{p}=1$. In this case we have an isometry $W_{p}=d_{Y}$. Under some further hypotheses, we can relate the functional $\mathcal{E}_{p}(f)$ to the metric measure foliation case.
@@
Let $(X,d)$ be a locally compact complete separable geodesic space, $Y$ a compact complete separable metric space, and $\pi:X\to Y$ a Borel map. Given $\mu\in\mathcal{M}_{+}(X)$, let $f_{\mu,\pi}$ be a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$. Suppose that the conditional measures $\mu_{y}$ have full support, i.e., $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$. Then, $\mathcal{E}_{p}(f_{\mu,\pi})=1$ if, and only if, $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$. Therefore, the energy-minimizing disintegration maps correspond to a metric measure foliation.
@@
[…]
$u(y^{\prime},y^{\prime\prime}):=\begin{cases}\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{d_{Y}(y^{\prime},y^{\prime\prime})}&~\text{,if}~~y^{\prime}\neq y^{\prime\prime}\\ 1&~\text{,if}~~y^{\prime}=y^{\prime\prime}\end{cases}$
Note that $u\geq 1$. By the hypothesis,
$\lim\limits_{\varepsilon\to 0}\sup\limits_{y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)}u(y^{\prime},y^{\prime\prime})=u(y,y)$
for every $y\in Y$. Then, $u$ is continuous at $(y,y)$ for every $y\in Y$. Moreover, $u$ is continuous everywhere else. In fact, since $\min_{f_{\mu,\pi}}\mathcal{E}_{p}(f_{\mu,\pi})=1$, we have $|\nabla f(y)|p=1$ for every $y\in Y$, and the disintegration map $f$ is continuous. Thus, it implies the continuity of $u$ everywhere else.
Since $Y$ compact, $Y\times Y$ is compact and by compactness, we have uniform continuity, that is, for all $\varepsilon>0$ there exists $\delta>0$ such that,
$d_{Y\times Y}((y,y^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$
implies $|u(y,y^{\prime})-u(\tilde{y},\tilde{y}^{\prime})|\leq\varepsilon$, for every ($y~)(y,y^{\prime}$),($\tilde$$\tilde{y}^{\prime}$), where
[…]
Note that $|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1$. Then, setting $y=\tilde{y}$, and if $d_{Y}((\tilde{y},\tilde{y}^{\prime}),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, we get $d_{Y\times Y}((y,y),(\tilde{y},\tilde{y}^{\prime}))\leq\delta$, Thus,
$|u(y,y)-u(\tilde{y},\tilde{y}^{\prime})|=u(\tilde{y},\tilde{y}^{\prime})-1\leq\varepsilon,$
[…]
$\sum_{k=0}^{n-1}d(x_{k},x_{k+1})=d(x,x^{\prime}).$
Then,
[…]
$\leq$ $\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(\pi^{-1}(\pi(x_{k})),\pi^{-1}(\pi(x_{k+1})))$
$\leq$ $(1+\varepsilon)\sum_{k=0}^{n-1}d(x_{k},x_{k+1})$
$=$ $(1+\varepsilon)d(x,x^{\prime})$
[…]
Let $\gamma$ be the optimal transport plan from $\mu_{y}$ to $\mu_{y^{\prime}}$, and consider, for all $(x,x^{\prime})\in\operatorname{supp}(\gamma)$, $\varphi(x)=d(x,\pi^{-1}(y^{\prime}))>d(x,x^{\prime})$. Then,
$\int\varphi(x)d\gamma=\int\varphi(x)d\mu_{y}=\int d(x,\pi^{-1}(y^{\prime}))d\mu_{y}>\int d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))d\mu_{y}$
that is a contradiction, since
$W_{1}(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}}):=\inf\limits_{\gamma\in\Pi(\mu_{\tilde{y}},\mu_{\tilde{y}^{\prime}})}\int d(x_{1},x_{2})~d\gamma(x_{1},x_{2}).$
An analogous argument applies to the case $p\in(1,\infty)$.
Remark 4.3.
The conclusion of Theorem A does not hold if full support is not assumed. In fact, consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is a ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, with major radius $y$. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ such that the disintegration of $\mu$ with respect to $\nu$ is $\mu_{y}=\delta_{(0,\frac{y}{\lambda})}$. In this case,
$d_{Y}(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
for every $y,y^{\prime}\in Y$. Therefore, we have isometry, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A graphical representation of the case $\lambda=1.5.$ is provided in Figure 2.
Figure 2. Illustration for the case in which the measures $\mu_{y}$ do not have full support: there is isometry between $W_{p}$ and $d_{Y}$, but $\{\pi^{-1}(y)\}$ (the ellipses) does not define a metric foliation.
Note that in the previous example, if $\lambda=1$, that is, the preimages are circles in $X$, then $\{\pi^{-1}(y)\}$ forms a metric foliation. In the context of Theorem A, we aim to demonstrate that the $p$-energy exhibits sensitivity to small perturbations of the foliation. To analyze this, we first evaluate the $1$-energy $\mathcal{E}_{1}$ for a metric measure foliation on $\mathbbm{R}^{2}$ with concentric circular leaves. We then generalize this analysis to cases where the leaves are slightly deformed into ellipses.

### v2

@@ 3. Preliminaries on disintegration maps
In this section, we discuss key properties of disintegration maps, as introduced in Definition 2.1 and developed in [PR25]. To start our discussion, suppose $X=Y=\mathbb{R}$, consider a measurable map $\pi:X\to Y$ and let $\gamma$ be a probability measure supported on the graph of $\pi$, i.e., $\gamma=(\mathrm{id},\pi)_{*}\mu$, where $\mu$ is some probability measure on $X$ and $(\mathrm{id},\pi)_{*}\mu$ is the push-forward of $\mu$ under the map $x\mapsto(x,\pi(x))$. Then, the disintegration of $\gamma$ with respect to $\mu$ is given by a family of Dirac measures $\gamma_{x}=\delta_{\pi(x)}$. Indeed, for every measurable set $A\subseteq X\times Y$, one has
@@
By definition of the disintegration, it holds
$\gamma(A)=\int_{X}\gamma_{x}(A_{x})\,d\mu(x)=\int_{X}\delta_{\pi(x)}(A_{x})\,d\mu(x),$
with $A_{x}=\{y\in Y:(x,y)\in A\}$. Then,
@@
In general, we aim to investigate under what conditions geometric information about $X$ can be derived via disintegration, specifically, through the properties of disintegration maps. A key starting point for this analysis is understanding the continuity of these maps. When referring to continuity of the disintegration map, we mean continuity with respect to weak convergence on $\mathscr{P}_{p}(X)$. In [PR25] some results are presented regarding the continuity of disintegration maps, namely:
@@
    [PR25, Propostion 5.2] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If a Borel map $\pi:X\to Y$ is such that $\nu:=\pi_{*}\mu$ is a Borel measure, then, the disintegration map $f:Y\to(\mathscr{P}_{p}(X),W_{p})$ of $\mu$ with respect to $\nu$ is nearly continuous, for $p\in[1,\infty)$. Here, a map $f$ on an metric space $(Y,\nu)$ is called nearly continuous if for each $\varepsilon>0$ there exist a closed subset $\mathcal{K}\subset Y$ with $\nu(Y\backslash\mathcal{K})<\varepsilon$ such that $f$ restricted to $\mathcal{K}$ is continuous.
@@
    [PR25, Propostion 5.4] Let $X$, $Y$ be locally compact complete separable metric spaces, and consider $\mu\in\mathcal{M}_{+}(X)$. If a Borel map $\pi:X\to Y$ is bijective and $\pi^{-1}$ is continuous, then the disintegration map $f$ of $\mu$ with respect to $\pi_{*}\mu$ is continuous.
@@
    [PR25, Propostion 5.8] Let $X$ be a locally compact complete separable metric space. Consider a metric foliation $\mathcal{F}$ of $X$, the quotient space $X^{*}$, and the quotient map $\pi:X\to X^{*}$. If $\mathcal{F}$ is a metric measure foliation, then the disintegration map of $\mu$ with respect to $\nu:=\pi_{*}\mu$ is continuous.
The continuity of disintegration maps enables us to relate paths in the chosen indexing metric space to corresponding paths of measures in Wasserstein space. For instance, in the case of the disintegration map $x\mapsto\delta_{x}$, there is a direct correspondence between a path in $X$ and the induced path of Dirac measures. In the context of a metric measure foliation, this relationship becomes more nuanced. Here, transversal paths (those crossing the leaves) must satisfy an additional condition concerning the distance between the supports of the conditional measures along the path, a consequence of the parallelism of the leaves.
In the next section we present an approach that allows us to study the relation between the relative position of the conditional measures in the Wasserstein space and the relative position of their respective supports, via disintegration map.
@@ 4. On foliations and disintegration maps
In this section, we will assume that the disintegration of measures holds for every point, rather than for $\nu$-almost every point, denoting the family as $\{\mu_{y}\}_{y\in Y}$ to explicitly indicate that the conditional measures are defined for every $y\in Y$. In fact, under some structural assumptions, the disintegration can hold for every $y\in Y$. This may occurs, for example, when $\pi$ is a trivial bundle with compact fiber, or when $Y$ is discrete. In such settings, $\mu_{y}$ can be defined point wise while preserving all disintegration properties.
Motivated by the concept of the metric derivative and the interplay between the relative positions of conditional measures and their supports, we introduce a notion of derivative for disintegration maps.
@@
Let $X$ and $Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}$ is a disintegration of $\mu$ with respect to $\nu=\pi_{*}\mu$. Let $f$ be the respective disintegration map. We write, for $y^{\prime},y^{\prime\prime}\in Y$,
$\rho(y^{\prime},y^{\prime\prime}):=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime})),$
and we define a notion of derivative of $f$ by
$|\nabla f(y)|_{p}:=\lim\limits_{\varepsilon\to 0}~\sup_{\begin{subarray}{c}\rho(y,y^{\prime})\leq\varepsilon\\ \rho(y,y^{\prime\prime})\leq\varepsilon\\ y^{\prime}\neq y^{\prime\prime}\end{subarray}}~\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{\rho(y^{\prime},y^{\prime\prime})}$ (4)
for $p\in[1,\infty)$, where $\mu_{y^{\prime}}=f(y^{\prime})$ and $\mu_{y^{\prime\prime}}=f(y^{\prime\prime})$ are conditional measures.
We warn the reader that $\rho$ may not satisfy all the distance axioms.
Note that $|\nabla f(y)|_{p}\geq 1$ everywhere. Indeed, since $\mu_{y^{\prime}}$ and $\mu_{y^{\prime\prime}}$ are supported on $\pi^{-1}(y^{\prime})$ and $\pi^{-1}(y^{\prime\prime})$, respectively, we have
@@
In the case of a $p$-metric measure foliation, the $p$-Wasserstein distance between conditional measures satisfies $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$, which implies $|\nabla f(y)|_{p}=1$ for every $y\in Y$. Meanwhile, $|\nabla f(y)|_{p}=1$ for every $y\in Y$ does not necessarily imply that the family of fibers $\{\pi^{-1}(y)\}$ is a $p$-metric measure foliation. However, this derivative can be used as a tool to characterize when $\{\pi^{-1}(y)\}$ forms a $p$-metric measure foliation, provided certain additional assumptions are made. Paralleling techniques from Geometric Analysis and Optimal Transport, we introduce a dense energy functional designed to further analyze the relationship between the distances of conditional measures and their supports.
Definition 4.2.
Let $X$ and $Y$ be locally compact complete separable metric spaces. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\nu=\pi_{*}\mu$. Let $f$ be the respective disintegration map. The $p$-energy of $f$ is given by
This energy definition involves analyzing $|\nabla f(y)|_{p}$ at every point $y\in Y$. Example 4.4 demonstrates a scenario in which the definition is satisfied for almost every point, yet the characterization we seek to prove does not hold. In the following theorem, we relate the value of this energy functional to settings in which the disintegration corresponds to a metric measure foliation case.
@@
Let $(X,d_{X})$ be a geodesic locally compact complete separable metric space. Consider a locally compact complete separable metric space $Y$, a measure $\mu\in\mathcal{M}_{+}(X)$ and a Borel map $\pi:X\to Y$ such that $\{\mu_{y}\}_{y\in Y}$ is a disintegration of $\mu$ with respect to $\nu:=\pi_{*}\mu$, with $\operatorname{supp}(\mu_{y})=\pi^{-1}(y)$. Let $f$ be the respective disintegration map. Then, $\mathcal{E}_{p}(f)=1$ if and only if $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$.
Remark 4.3.
The assumptions on the space $Y$ are made only to guarantee the existence of the disintegration of $\mu$ with respect to $\pi_{*}\mu$, according [PR25, Theorem A]. Indeed, the claim of our theorem only concerns the distance between the fibers, and not the original distance on $Y$.
@@
The demonstration will be carried out in three steps. Step 1 shows that $p$-metric measure foliation implies $\mathcal{E}_{p}(f)=1$, and Step 2 and Step 3 show that if $\mathcal{E}_{p}(f)=1$, then $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$. More specifically, Step 2 shows that the Wasserstein distance between two conditional measures coincides with the distance between the corresponding fibers; Step 3 shows the metric foliation property, that is, all points on a fiber have the same distance to another fiber.
In the sequel, we consider the distance between the preimages of $y^{\prime}$ and $y^{\prime\prime}$ to be given by $\rho(y^{\prime},y^{\prime\prime})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))=\inf\{d_{X}(x^{\prime},x^{\prime\prime}):\mbox{$x^{\prime}\in\pi^{-1}(y^{\prime})$}$, $x^{\prime\prime}\in\pi^{-1}(y^{\prime\prime})\}$.
Step 1. Suppose that $\{\pi^{-1}(y)\}$ defines a $p$-metric measure foliation on $X$. Then, for any $y^{\prime},y^{\prime\prime}\in Y$, holds $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$. Therefore, for any $y\in Y$, one has
$|\nabla f(y)|_{p}=\lim\limits_{\varepsilon\to 0}~\sup_{\begin{subarray}{c}\rho(y,y^{\prime})\leq\varepsilon\\ \rho(y,y^{\prime\prime})\leq\varepsilon\\ y^{\prime}\neq y^{\prime\prime}\end{subarray}}~\frac{W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})}{\rho(y^{\prime},y^{\prime\prime})}=1,$
which implies $\mathcal{E}_{p}(f)=1$.
Step 2. Suppose $\mathcal{E}_{p}(f)=1$. We want to show that in this case, for any $y^{\prime},y^{\prime\prime}\in Y$, it holds that $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$.
Consider $\varepsilon>0$ and $y_{0},y_{1}\in Y$. Then there exist $x_{0}\in\pi^{-1}(y_{0})$, $x_{1}\in\pi^{-1}(y_{1})$, such that $d_{X}(x_{0},x_{1})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))+\varepsilon$. Let $\gamma:[0,1]\to X$ be a constant speed minimizing geodesic such that $\gamma(0)=x_{0}$ and $\gamma(1)=x_{1}$. Define $y_{t}:=\pi(\gamma(t))$, $t\in[0,1]$. Note that, for every $s,t\in[0,1]$, it holds
$W_{p}(\mu_{y_{s}},\mu_{y_{t}})\geq d(\pi^{-1}(y_{s}),\pi^{-1}(y_{t}))$ (5)
and
$d(\pi^{-1}(y_{s}),\pi^{-1}(y_{t}))\leq|s-t|d_{X}(x_{0},x_{1}).$ (6)
Define $\varphi(t):=W_{p}(\mu_{y_{t}},\mu_{y_{0}})$. We will show that $\varphi$ is continuous and that
$\partial_{t}^{+}(\varphi(t)):=\limsup_{h\to 0}\frac{\varphi(t+h)-\varphi(t)}{h}\leq d_{X}(x_{0},x_{1}).$
Indeed,
$\limsup_{h\to 0}\left|\frac{\varphi(t+h)-\varphi(t)}{h}\right|=$ $\limsup_{h\to 0}\left|\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{0}})-W_{p}(\mu_{y_{t}},\mu_{y_{0}})}{h}\right|$
$\leq$ $\limsup_{h\to 0}\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{t}})}{|h|}$
$\leq$ $d_{X}(x_{0},x_{1})\limsup_{h\to 0}\frac{W_{p}(\mu_{y_{t+h}},\mu_{y_{t}})}{d(\pi^{-1}(y_{t+h}),\pi^{-1}(y_{t}))}$
$\leq$ $d_{X}(x_{0},x_{1})|\nabla f(y_{t})|_{p}$
$\leq$ $d_{X}(x_{0},x_{1})\mathcal{E}_{p}(f)$
$=$ $d_{X}(x_{0},x_{1}).$
Here, the first inequality follows from the triangle inequality for the Wasserstein distance; the second one follows from (6); the third one follows from the definition of the gradient (4); the last equation follows from the assumption $\mathcal{E}_{p}(f)=1$. Now, the continuity of $\varphi$ follows easily.
By [HT06, Theorem 2], we can integrate the inequality $\partial_{t}^{+}(\varphi(t))\leq d_{X}(x_{0},x_{1})$, giving $\varphi(1)-\varphi(0)\leq d_{X}(x_{0},x_{1})$. Thus,
$W_{p}(\mu_{y_{0}},\mu_{y_{1}})\leq d_{X}(x_{0},x_{1})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))+\varepsilon.$
Since $\varepsilon$ is arbitrary, taking the limit $\varepsilon\to 0$, we have
$W_{p}(\mu_{y_{0}},\mu_{y_{1}})\leq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1})).$ (7)
By (5) and (7) we obtain $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))$.
Step 3. Keep assuming $\mathcal{E}_{p}(f)=1$. We will show that for every $y_{0},y_{1}\in Y$, and for every $x_{0}\in\pi^{-1}(y_{0})$, it holds $d(x_{0},\pi^{-1}(y_{1}))=d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$.
Note that, by definition, one has $d(x_{0},\pi^{-1}(y_{1}))\geq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$.
Consider a function $r:\pi^{-1}(y_{0})\to\mathbb{R}^{+}_{0}$ given by $r(x):=d(x,\pi^{-1}(y_{1}))$. Note that $r(x)\geq d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$ for every $x\in\pi^{-1}(y_{0})$ and $r$ is continuous. Moreover, note that for $(x,x^{\prime})\in\pi^{-1}(y_{0})\times\pi^{-1}(y_{1})$, one has $d_{X}(x,x^{\prime})\geq r(x)$. Let $\eta$ be an optimal transport plan from $\mu_{y_{0}}$ to $\mu_{y_{1}}$. Then, by Step 2, it holds
$\int d_{X}(x,x^{\prime})^{p}d\eta=W^{p}_{p}(\mu_{y_{0}},\mu_{y_{1}})=[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}.$
Furthermore, one has
$\int d_{X}(x,x^{\prime})^{p}d\eta\geq\int r(x)^{p}d\mu_{y_{0}}\geq[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}.$
Since $r^{p}$ is continuous and $\operatorname{supp}(\mu_{y_{0}})=\pi^{-1}(y_{0})$, for every $x\in\pi^{-1}(y_{0})$ we have $r(x)^{p}=[d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))]^{p}$ . Therefore, for every $x_{0}\in\pi^{-1}(y_{0})$, we have $d(x_{0},\pi^{-1}(y_{1}))=d(\pi^{-1}(y_{0}),\pi^{-1}(y_{1}))$
Step 2 and Step 3 complete the proof.
In Theorem A, we characterize a foliation on $X$ via the energy of a disintegration map. We emphasize that the hypotheses of conditional measures having full support and the analysis of the energy functional over all points are fundamental to the result. In the following examples, we present cases that justify the use of such assumptions.